Project

General

Profile

Actions

---------------------------------------------------
List of HVs:
board1, 6kV, 8-ch:
Name:_GEM00,_GEM01,_GEM10,_GEM11,__,__,__,_Cer_D
HV:_______0,_____0,_____0,_____0,_0,_0,_0,__1000

board2, 2kV, 16-ch:
Name:_SC_D,_SC_A,_SC_B,_X,_LASPD_t,_LASPD_b,_SC_C,_PreSh_L(SDU#1),_PreSh_R(THU#1),_PreSh_T(SDU#2),_Shower_L(SDU#4),_Shower_R(THU#2),_Shower_T(SDU#5),__CerA,__CerB,__CerC
HV:___1200,_1600,_1650,_0,_____850,____1000,_1600,___________1515,___________1800,___________1730,_____________930,_____________950,_____________850,__1000,__1000,__1000


3944 By Jixie: debug FADC trigger

3949 cosmic(6h) and tune beam no target(3h). By Jixie: debug FADC trigger

3950 Pedestal. No HV
3951 cosmic, no beam. debug FADC trigger
3952 debug gem ssp
3953 cosmic, no beam. debug FADC trigger
3954 production. 21:16 Jan 11st - 23:14 Jan 11st
daq rate = 12.5kHz

3955 production. 22:16 Jan 11st - 23:32 Jan 11st
daq rate = 12kHz

change prescale factors, disable internal pulse trigger for the first 100k events /*Set prescale for each TS#*/
tiSetInputPrescale(1,0);
tiSetInputPrescale(2,0);
tiSetInputPrescale(3,5);
tiSetInputPrescale(4,3);
tiSetInputPrescale(5,3);
tiSetInputPrescale(6,0); use hallc_fadc_ssp config

3956 trash. 22:34 Jan 11st
tried to use hallc_fadc_ssp config, could not start CODA

use hallc_fadc config

3957 production. 23:37 Jan 11st - 23:45 Jan 11st
daq rate = 12.5kHz @ 2uA , this is the maximum of the daq rate for uvasoild2
Threshold: Cer_Sum(30mV), Shower_Sum(10mV),SC_D(Kedi6-Tile,10mv),SC_B(downstream of Shower,10mV)

3958 production. 23:48 Jan 11st - 00:47 Jan 12th
Threshold: Cer_Sum(100mV), Shower_Sum(10mV),SC_D(Kedi6-Tile,10mv),SC_B(downstream of Shower,10mV)

Turn off HVs

Jan 12th, 12:30

3959-3971 Trash, debug daq

-----------------------------------------------------------------------
Turn on HVs, all Cer HVs are changed from 1000V to 950V, Shower_T(SDU#5) lower from 850V to 800V

---------------------------------------------------
List of HVs:
board1, 6kV, 8-ch:
Name:_GEM00,_GEM01,_GEM10,_GEM11,__,__,__,_Cer_D
HV:_______0,_____0,_____0,_____0,_0,_0,_0,__950

board2, 2kV, 16-ch:
Name:_SC_D,_SC_A,_SC_B,_X,_LASPD_t,_LASPD_b,_SC_C,_PreSh_L(SDU#1),_PreSh_R(THU#1),_PreSh_T(SDU#2),_Shower_L(SDU#4),_Shower_R(THU#2),_Shower_T(SDU#5),__CerA,__CerB,__CerC
HV:___1200,_1600,_1650,_0,_____850,____1000,_1600,___________1515,___________1800,___________1730,_____________930,_____________950,_____________800,___950,___950,___950


only TS#6 is enabled, PS#(15,15,15,15,15,15,0)
Threshold: Cer_Sum(300mV), Shower_Sum(10mV),SC_D(Kedi6-Tile,10mv),SC_B(downstream of Shower,10mV)

Cer_HV = 950V
3972 pulse, no beam. 18:11 Jan 12th - 18:18 Jan 12th, Cer_HV = 950V
3973 pulse, no beam. 18:19 Jan 12th - 18:24 Jan 12th, Cer_HV = 950V
15.5kHz @ no beam, pulse trigger

Cer_HV = 970V
3974 pulse, no beam. 18:26 Jan 12th - 18:32 Jan 12th, Cer_HV = 970V
3975 pulse, no beam. 18:32 Jan 12th - 18:52 Jan 12th, Cer_HV = 970V

Cer_HV = 930V
3976 pulse, no beam. 18:54 Jan 12th - 19:01 Jan 12th, Cer_HV = 930V
3977 pulse, no beam. 19:03 Jan 12th - 19:09 Jan 12th, Cer_HV = 930V

Cer_HV = 1000V
3978 pulse, no beam. 19:54 Jan 12th - 20:06 Jan 12th, Cer_HV = 1000V
3979 pulse, no beam. 20:06 Jan 12th - 20:16 Jan 12th, Cer_HV = 1000V

Cer_HV = 1050V
3980 pulse, no beam. 20:17 Jan 12th - 20:24 Jan 12th, Cer_HV = 1050V
3981 pulse, no beam. 20:25 Jan 12th - 20:39 Jan 12th, Cer_HV = 1050V
-----------------------------------------------------------------------

3982 cosmic production,
first 100k is pulse trigger, 22:36 - 23:28 Jan 12th
debug TS#6, all trigger bits enabled, PS#(0,0,0,0,0,0,0)
Threshold: Cer_Sum(300mV), Shower_Sum(10mV),SC_D(Kedi6-Tile,10mv),SC_B(downstream of Shower,10mV)

cosmic test for Shower only, HVs for others are off, first 100k is pulse trigger
TS#4 only PS#(15,15,15,0,15,15)
Threshold: Cer_Sum(300mV), Shower_Sum(10mV),SC_D(Kedi6-Tile,10mv),SC_B(downstream of Shower,10mV)

3983 cosmic production, 23:34 - 00:25 Jan 13th
Name:_Shower_L(SDU#4),_Shower_R(THU#2),_Shower_T(SDU#5)
HV:_______________930,_____________950,_____________800

00:27 Jan 13th, turn all HVs off

3984 cosmic production, 08:25 Jan 13th - 13:03 Jan 13th
Name:_Shower_L(SDU#4),_Shower_R(THU#2),_Shower_T(SDU#5)
HV:______________1030,____________1050,_____________850

3985 cosmic production, 13:04 Jan 13th - 17:13 Jan 13th
Name:_Shower_L(SDU#4),_Shower_R(THU#2),_Shower_T(SDU#5)
HV:______________1030,____________1050,_____________850
see 12 minutes of beam ... detectors are fully saturated

3986 cosmic production, 17:13 Jan 13th - 18:32 Jan 13th
Name:_Shower_L(SDU#4),_Shower_R(THU#2),_Shower_T(SDU#5)
HV:______________1030,____________1050,_____________850
see 5 minutes of beam ... detectors are fully saturated

all HVs are off
all 6 trigger bits are on, no prescale factors dark current run (beam on but HV off, pulse trigger)
3987 cosmic production, 18:36 Jan 13th - 18:42 Jan 13th
I tried to get beam on data, but beam is off

3988 dark current run, 11:05 Jan 14th - 11:11 Jan 14th
beam current 40uA, HV off, pulse trigger

3989 dark current run, 11:12 Jan 14th - 11:20 Jan 14th
beam current 40uA, HV off, pulse trigger

=======================================================================
5uA to test the prescale factors
HV for CerD is off (forgot to turn on)

Name:_SC_D,_SC_A,_SC_B,_X,_LASPD_t,_LASPD_b,_SC_C,_PreSh_L(SDU#1),_PreSh_R(THU#1),_PreSh_T(SDU#2),_Shower_L(SDU#4),_Shower_R(THU#2),_Shower_T(SDU#5),__CerA,__CerB,__CerC
HV:___1200,_1600,_1650,_0,_____850,____1000,_1600,___________1515,___________1800,___________1730,____________1030,_____________1050,____________850,___950,___950,___950

3990 all PS=0 TS1-6 daq rate=15.5kHz
3991 all PS=0 TS1-6 daq rate=15.5kHz
3992 PS4=5, TS1-6 daq rate=15.5kHz
3993 TS4 only PS4=5 daq rate=12.5kHz

=======================================================================
Increase shower_T HV from 850 to 900
Decrease shower_R HV from 1050 to 1030
HVs for SC_D, SC_A, LASPD_T,LASPD_B, PreSh_# were all got lower

Name:_SC_D,_SC_A,_SC_B,_X,_LASPD_t,_LASPD_b,_SC_C,_PreSh_L(SDU#1),_PreSh_R(THU#1),_PreSh_T(SDU#2),_Shower_L(SDU#4),_Shower_R(THU#2),_Shower_T(SDU#5),__CerA,__CerB,__CerC
HV:___1150,_1550,_1650,_0,_____800,_____900,_1600,___________1470,___________1740,___________1680,____________1030,_____________1030,____________900,___930,___930,___930

CerD HV is also 930V

3994-3995 trash, HV boardB tripped

3996 dark current run, beam off, HVs all on, pulse trigger, 930V
3997 dark current run, beam off, HVs all on, pulse trigger, 930V

3998 dark current run, beam off, HVs all on, pulse trigger, 940V
3999 dark current run, beam off, HVs all on, pulse trigger, 940V

4000 dark current run, beam off, HVs all on, pulse trigger, 920V
4001 dark current run, beam off, HVs all on, pulse trigger, 920V

4002 dark current run, beam off, HVs all on, pulse trigger, 910V
4003 dark current run, beam off, HVs all on, pulse trigger, 910V

4004 dark current run, beam off, HVs all on, pulse trigger, 950V
4005 dark current run, beam off, HVs all on, pulse trigger, 950V

4006 dark current run, beam off, HVs all on, pulse trigger, 900V
4007 dark current run, beam off, HVs all on, pulse trigger, 900V

4008 dark current run, beam off, HVs all on, pulse trigger, 960V
4009 dark current run, beam off, HVs all on, pulse trigger, 960V

====================================================================
Cer HVs go back to 950V, 1uA CW beam on He3 target

4010 beam on, but only pulse trigger
4011 first 100k is pulse trigger, 50Hz rate without beam. TS1 to 6 are all on, PS4=5; NO real beam on this run
4012 disabled pulse trigger. All TS#. PS4=5, no beam
4013 disabled pulse trigger. All TS#. PS4=9, daq rate = 13.5kHz,1uA beam
4014 disabled pulse trigger. All TS#. PS4=11, daq rate = 13.5kHz,1uA beam
4015 disabled pulse trigger. All TS#. PS4=15, PS5=15, daq rate = 13.5kHz,1uA beam
4016 disabled pulse trigger. only TS1,TS2,TS3,TS6 enabled, PS1=PS2=PS3=PS6=0, daq rate = 13.5kHz, 1uA beam
4017 disabled pulse trigger. only TS1,TS2,TS3,TS6 enabled, PS1=PS3=PS6=15, PS2=5, no beam
4018 disabled pulse trigger. only TS1,TS2,TS3,TS6 enabled, PS1=PS3=PS6=15, PS2=5, 2uA beam. 12.8KHz
4019 disabled pulse trigger. only TS1,TS2,TS3,TS6 enabled, PS1=PS3=PS6=15, PS2=10, 2uA beam. 8.5KHz
4020 disabled pulse trigger. only TS1,TS2,TS3,TS6 enabled, PS1=PS3=PS6=15, PS2=12, 2uA beam. 2.7KHz
4021 disabled pulse trigger. only TS1,TS2,TS3,TS6 enabled, PS3=PS6=15, PS1=0,PS2=12, 2uA beam. 11KHz
4022 disabled pulse trigger. only TS1,TS2,TS3,TS6 enabled, PS3=PS6=15, PS1=5,PS2=12, 1.8uA beam. 5.7KHz

Conclusion: on He3 target, PS1=5 will have 1.8kHz/uA for TS#1, PS2=12 will have 1.4kHz/uA for TS#2.
We still need to determine PS for PS4 and PS5.
After analyze run 4022, I found that for PS3=15, TS#3 has 28Hz/2uA. For PS6=15, TS#6 has 8Hz/2uA. Scaling PS from
15 to 10, we got TS#3 rate as 450Hz/uA and TS#6 rate 130Hz/uA.

===================================================
20230120 pedestal runs, beam off, pulse trigger, HV scan

Name:_SC_D,_SC_A,_SC_B,_X,_LASPD_t,_LASPD_b,_SC_C,_PreSh_L(SDU#1),_PreSh_R(THU#1),_PreSh_T(SDU#2),_Shower_L(SDU#4),_Shower_R(THU#2),_Shower_T(SDU#5),__CerA,__CerB,__CerC
HV:___1150,_1550,_1650,_0,_____800,_____900,_1600,___________1470,___________1740,___________1680,____________1030,_____________1030,____________900,___930,___930,___930

For Cherenkov: ch13,ch14,ch15, E08C2_ch07: 910, 930, 950, 970
For Shower: ch10,ch11: 900, 950, 1000, 1050
For Shower: ch12: 800, 850, 900, 950

ch 10   11   12   13   14  15    07
4023, 4024: HV(ch10-15,E08C2_ch07) : 800, 900, 900, 910, 910, 910, 910
4025, 4026: 950, 950, 850, 930, 930, 930, 930
4027, 4028: 1000,1000, 900, 950, 950, 950, 950
4029, 4030: 1050,1050, 950, 970, 970, 970, 970
4031, 4032: 1030,1030, 900, 950, 950, 950, 950

===================================================

Name:_SC_D,_SC_A,_SC_B,_X,_LASPD_t,_LASPD_b,_SC_C,_PreSh_L(SDU#1),_PreSh_R(THU#1),_PreSh_T(SDU#2),_Shower_L(SDU#4),_Shower_R(THU#2),_Shower_T(SDU#5),__CerA,__CerB,__CerC
HV:___1150,_1550,_1650,_0,_____800,_____900,_1600,___________1470,___________1740,___________1680,____________1030,_____________1030,____________900,___930,___930,___930

turn on FADC threshold to remove no pulse events
slot 7: FADC250_ALLCH_TET 31 31 31 31 31 31 31 31 31 31 31 31 31 31 31 31
slot 8: FADC250_ALLCH_TET 10 10 10 20 10 10 10 5 5 5 5 5 5 15 15 25

4033 debug FADC config, no beam, HV on. 5M pulse trigger
4034 debug FADC config, no beam, HV on. 5M pulse trigger

turn on slot 8 FADC threshold to remove no pulse events
slot 7: FADC250_ALLCH_TET 31 31 31 31 31 31 31 31 31 31 31 31 31 31 31 31
slot 8: FADC250_ALLCH_TET 10 10 10 20 10 10 10 10 10 10 10 10 10 20 20 25
4035 trash
4036 debug FADC config, no beam, HV on. 5M pulse trigger

==============
4037 disabled pulse trigger, cosmic only, beam off HV on

==============
4038-4040 trash, check daq, adjust threshold

========================================================
20230123 threshold
Threshold is applied to the height of the waveform:
TS1: CerSum(200mv or 410 adc unit) & ShowerSum(70mv or 143 adc unit);
TS2: SC_D(20mv or 164 adc unit) & ShowerSum(70mv or 143 adc unit) & SC_B(15mv or 123 adc unit).
TS3: 2 out of these 3 inputs: CerSum(200mv or 410 adc unit), ShowerSum(50mv or 103 adc unit), SC_D(20mv or 164 adc unit);
TS4: ShowerSum(100mv or 205 adc unit)
TS5: SC_B(20mv or 164 adc unit).
TS6: CerA#,CerB#,CerC#,CerD#(15mv or 31 adc unit) for each Cherenkov individual input

5uA tune beam on LHe4,
4041 5uA tune beam on LHe4, PS=10,10,10,10,10,10, rate=250Hz
4042 5uA tune beam on LHe4, PS=4,4,3,1,0,0, rate=2.4kHz
4043 5uA tune beam on LHe4, PS=3,3,2,0,0,0, rate=4kHz

2uA CW beam on LHe4,
4044 all TS1-6; PS=7,7,5,5,3,0, 11.5kHz, daq saturated
4045 all TS1-6; PS=8,8,5,3,4,0, 11.2kHz, daq saturated
4046 all TS1-6; PS=8,8,5,3,3,1, 11.2kHz, daq saturated
4047 all TS1-6; PS=9,9,5,3,3,0, 11.2kHz, daq saturated

2uA CW beanm on LHe4,
4048 TS#4 only, PS4=3, rate=10.0kHz
4049-4050 trash
4051 TS#4 only, PS4=4, rate=5.9kHz
4052-4054 trash, TS#5 only, no rate, TS#5 cable connection might have problem
4055 TS#4 only, PS4=4, rate=5.6kHz
turn off sw test pulse in diaggui, this test inserts 100Hz pulse into FADC
do not know how it affects the above data
4056 trash, TS#6 only, no rate,
4057 TS#3 only, PS3=5, rate=2.0kHz
4058 TS#3 only, PS3=3, rate=6.8kHz
4059 trash
4060 TS#2 only, PS2=10, rate=1.1kHz
4061 TS#2 only, PS2= 7, rate=7.6kHz
4062 TS#1 only, PS1= 9, rate=2.8kHz
4063 TS#1 only, PS1= 7, rate=9.2kHz
4064 All TS, PS=9,9,5,6,6,1, rate=7.8kHz,

4065 All TS, PS=9,9,5,6,6,1, 20uA beam on LHe4, rate=11.2kHz, daq saturated

========================================================
4066 pedestal verification, beam off, HV off, pulse trigger, 500k

------------------------------------------------------------------------
ToFix:
Jimmy reported that the level1 tree has bug that some variables (cer16) are still
keeping the value of previous event if there is no hit in current event. (run4015)
------------------------------------------------------------------------

Mistakes for runs from 3972 to 4066
1. The vmeDSC has positive pulse output if one follows the +|- label. NIM logic board expects
negative pulse inputs. Therefore our TS#1,TS#2 and TS#3 in run 3972-4066 are not what we designed.
TS#1 and TS#2 are logical "OR". TS#3 is complicated to express.

Mistakes for runs between 4023 and 4066:
1. TS#4 is actually triggered by SC_B due to cable wrongly shift by 1 pin.
On Jan 20 I tried to move TS#4(vmeDSC_ch12) and TS#5(vmeDSC_ch11) to vmeDSC_ch14 and vmeDSC_ch15.
But they are actually shifted by one pin location, which are vmeDSC_ch15 and vmeDSC_ch16.
Due to this mistake, trigger-bit 4 (0x1000) is actually triggered by SC_B (the SC behind shower),
Runs 4023-4066 are affected.

=========================================================
20230124
Fixed ECL cable invertly connection problem, FADC TrigOut is connected to vmeDSC_ch16, then connecting to TS#6
Verified TS#1-5 using function generator, still need to verify TS#6 from data

4067 all TS1-6, no PS, pulse trigger, no beam
4068 all TS1-6, no PS, pulse trigger, no beam
4069 all TS1-6, no PS, cosmic trigger trigger, no beam
cosmic trigger threshold
Threshold is applied to the height of the waveform:
TS1: CerSum(150mv or 307 adc unit) & ShowerSum(30mv or 61 adc unit);
TS2: SC_D(20mv or 164 adc unit) & ShowerSum(30mv or 61 adc unit) & SC_B(15mv or 123 adc unit).
TS3: 2 out of these 3 inputs: CerSum(150mv or 307 adc unit), ShowerSum(30mv or 61 adc unit), SC_D(20mv or 164 adc unit);
TS4: ShowerSum(50mv or 102 adc unit).
TS5: SC_B(20mv or 164 adc unit).
TS6: CerA#,CerB#,CerC#,CerD#(15mv or 31 adc unit) for each Cherenkov individual input

=========================================================

to do:
look into run 4055 for Shower_R and SC_A

Raise TS4 threshold to 700mV and take some runs.
For TS4-only run, please adjust PS to keep rate at 3kHz. (If PS is too high, raise the threshold to 800mV or 900mV or 1V).
Need dedicated runs for each of TS4,5,1,2,3, in that order, and ideally also one run with all triggers on.

4071 40uA beam on Carbon target, TS#4 only, PS=0, only Shower HVs are on. ShowerSum threshold is 85mV, daq rate is 4.5kHz.
4072 40uA beam on Carbon target, TS#5 only, PS=0, only SC_B HV is on. SC_B threshold is 90mV, daq rate is 4.8kHz.
4073 trash
4074 40uA beam on Carbon target, TS#2 only, PS2=5, all HVs are on. threshold: SC_B(15mV),SC_D(20mV),ShowerSum(50mV), daq rate is 7.7kHz.
4075 trash
4076 40uA beam on Carbon target, TS#3 only, PS3=10, all HVs are on. threshold: CerSum(200mV),SC_D(20mV),ShowerSum(50mV), daq rate is 5.5kHz.
4077 40uA beam on Carbon target, TS#1 only, PS1=7, all HVs are on. threshold: CerSum(200mV),ShowerSum(70mV), daq rate is 0.7kHz.
4078 40uA beam on Carbon target, TS#1 only, PS1=4, all HVs are on. threshold: CerSum(200mV),ShowerSum(70mV), daq rate is 4.8kHz.
4079 40uA beam on Carbon target, TS#6 only, PS6=5, all HVs are on. threshold: CerSum(200mV), daq rate is 0Hz. MaPMTs are fully saturated.
4080 40uA beam on Carbon target, TS#5 only, PS=8, only SC_B HV is on. SC_B threshold is 30mV, daq rate is 6.8kHz.

I doubt the signal cable of left and right shower are swapped
4081 debug daq, only left preshower and left shower on, TS4 only, PS4=1, threshold=85mV
4082 debug daq, only right preshower and right shower on, TS4 only, PS4=1, threshold=85mV
4083 debug daq, only top preshower and top shower on, TS4 only, PS4=1, threshold=85mV

Result(conclusion): cable connections are all good.

=========================================================

A) What type of input pulse TI expects?
The TI we are using expects a positive differential pulse (meaning positive pulse on the + pin,
and simultaneously a negative pulse on the - pin).
The output of most level translator used in JLab can connect directly to TI, but not any logic unit output.
Its manual can be found here:
https://coda.jlab.org/drupal/system/files/pdfs/HardwareManual/TI/TI.pdf

B) can we tell readout to read FADC waveform even though it can not pass the given threshold?

JLab FADC250 can force all channels to be readout even with a threshold still specified.
The flag/mask, FADC250_TET_IGNORE_MASK, tells the readout whether or not to use the threshold (but the trigger will).
For example:
FADC250_TET_IGNORE_MASK 1 1 1 1 1 1 1 1 1 1 1 1 1 1 1 1

to do:
1) Addjust the following thresholds
Trig5: SC_B 10mV
Trig4: addjust threshold to get 3-4 kHz rate when PS4=0;
Trig2: SC_B 10mV, SC_D 10mV, ShowerSum 10mV

**Trig1 and Trag3 not mentioned in the meeting, should I change ShowerSum for these 2???

2)take dedicated runs for each trigger type in this order : TS4, TS5, TS1, TS2, TS3, TS6
It will take about 45s-60s to change TS configuration.
Each run will take about 300k to 400k triggers for 60s of beam time.
in total 12 minutes.

3)take one run with all triggers on. (2 minutes including TS configuration changing time)

===================================================================
20230127

4084 pedestal run. No beam HVs on, Preshower HVs stayed at ____1470,__________1740,___________1680,
4085 pedestal run. No beam HVs on, Preshower HVs changed back to __1515,_______1800,___________1730,
4086 pedestal run. No beam HVs off

15:11 all HVs on
---------------------------------------------------
List of HVs:
board1, 6kV, 8-ch:
Name:_GEM00,_GEM01,_GEM10,_GEM11,__,__,__,_Cer_D
HV:_______0,_____0,_____0,_____0,_0,_0,_0,___950

Name:_SC_D,_SC_A,_SC_B,_X,_LASPD_t,_LASPD_b,_SC_C,_PreSh_L(SDU#1),_PreSh_R(THU#1),_PreSh_T(SDU#2),_Shower_L(SDU#4),_Shower_R(THU#2),_Shower_T(SDU#5),__CerA,__CerB,__CerC
HV:___1150,_1550,_1650,_0,_____800,_____900,_1600,___________1470,___________1740,___________1680,____________1030,_____________1030,____________900,___950,___950,___950

20230127 threshold
Threshold is applied to the height of the waveform:
TS1: CerSum(200mv or 410 adc unit) & ShowerSum(70mv or 143 adc unit);
TS2: SC_D(10mv or 82 adc unit) & ShowerSum(10mv or 21 adc unit) & SC_B(10mv or 82 adc unit).
TS3: 2 out of these 3 inputs: CerSum(200mv or 410 adc unit), ShowerSum(50mv or 103 adc unit), SC_D(10mv or 82 adc unit);
TS4: ShowerSum(200mv or 410 adc unit)
TS5: SC_B(10mv or 82 adc unit).
TS6: CerA#,CerB#,CerC#,CerD#(15mv or 31 adc unit) for each Cherenkov individual input

2uA cw beam on He3 target

4087 TS#4 only, PS4=0, threshold=67mV, daq rate=12kHz, saturated
4088 TS#4 only, PS4=0, threshold=70mV, daq rate=12kHz, saturated
4089 TS#4 only, PS4=0, threshold=75mV, daq rate=12kHz, saturated
4090 TS#4 only, PS4=0, threshold=80mV, daq rate=12kHz, saturated
4091 TS#4 only, PS4=0, threshold=100mV, daq rate=12kHz, saturated
4092 TS#4 only, PS4=0, threshold=200mV, daq rate=12kHz, saturated
4093 TS#4 only, PS4=5, threshold=200mV, daq rate=1.2kHz
4094 TS#4 only, PS4=3, threshold=200mV, daq rate=5kHz
4095 TS#2 only, PS2=10,threshold=SC_D(10mV),SC_B(10mV),ShowerSum(10mV), daq rate=5Hz
4096 TS#2 only, PS2=3, threshold=SC_D(10mV),SC_B(10mV),ShowerSum(10mV), daq rate=550Hz
4097 TS#2 only, PS2=0, threshold=SC_D(10mV),SC_B(10mV),ShowerSum(10mV), daq rate=2.5kHz
4098 TS#1 only, PS1=4, threshold=CerSum(200mV),ShowerSum(50mV), daq rate=3.2kHz
4099 TS#1 only, PS1=2, threshold=CerSum(200mV),ShowerSum(50mV), daq rate=9.5kHz
4100 TS#1 only, PS1=4, threshold=CerSum(200mV),ShowerSum(50mV), daq rate=3.2kHz
4101 TS#1-6 all, PS1-6=4,0,10,4,10,5, daq rate=5.2kHz

===============================================================================
4102-4109 debug GEM daq

---------------------------------------------------
List of HVs:
board1, 6kV, 8-ch:
Name:_GEM00,_GEM01,_GEM10,_GEM11,__,__,__,_Cer_D
HV:_______0,_____0,_____0,_____0,_0,_0,_0,___950

Name:_SC_D,_SC_A,_SC_B,_X,_LASPD_t,_LASPD_b,_SC_C,_PreSh_L(SDU#1),_PreSh_R(THU#1),_PreSh_T(SDU#2),_Shower_L(SDU#4),_Shower_R(THU#2),_Shower_T(SDU#5),__CerA,__CerB,__CerC
HV:___1150,_1550,_1650,_0,_____800,_____900,_1600,___________1470,___________1740,___________1680,_____________1030,____________1030,____________900,___950,___950,___950

20230201
1. Chagned GEM gas bottle
2. Connected GEM cables
3. Flipped polarity for the inputs of TS4,TS5,TS6

4110 pedestal run, beam off, HVs off, pulse trigger
4111 pedestal run, beam off, HVs off, pulse trigger
4112 pedestal run, beam off, HVs on, pulse trigger
4113 pedestal run, beam off, HVs on, pulse trigger

4114-4118 GEM pedestal run

20230203

---------------------------------------------------
List of HVs:
board1, 6kV, 8-ch:
Name:_GEM00,_GEM01,_GEM10,_GEM11,__,__,__,_Cer_D
HV:_______0,_____0,_____0,_____0,_0,_0,_0,___950

Name:_SC_D,_SC_A,_SC_B,_X,_LASPD_t,_LASPD_b,_SC_C,_PreSh_L(SDU#1),_PreSh_R(THU#1),_PreSh_T(SDU#2),_Shower_L(SDU#4),_Shower_R(THU#2),_Shower_T(SDU#5),__CerA,__CerB,__CerC
HV:___1150,_1550,_1650,_0,_____800,_____900,_1600,___________1470,___________1740,___________1680,_____________900,____________1030,____________1030,___950,___950,___950

20230203 threshold
Threshold is applied to the height of the waveform:
TS1: CerSum(100mv or 205 adc unit) & ShowerSum(150mv or 307 adc unit);
TS2: SC_D(5mv or 41 adc unit) & ShowerSum(5mv or 11 adc unit) & SC_B(6mv or 49 adc unit).
TS3: disabled
TS4: ShowerSum(300mv or 615 adc unit)
TS5: SC_B(6mv or 49 adc unit).
TS6: CerA#,CerB#,CerC#,CerD#(15mv or 31 adc unit) for each Cherenkov individual input

4019 GEM pedestal run, HV off, 40uA beam on He3 target

2uA CW beam on He3 target
all HVs are on without GEM, hallc_fadc config

4120 TS4 only, PS4=3, rate<600Hz
4121 TS4 only, PS4=0, rate=1.2kHz
4122 TS2 only, PS2=0, rate=5.4kHz
4123 TS1 only, PS1=5, rate<25Hz
4124 TS1 only, PS1=0, rate=600Hz
4125 TS1,TS2,TS4,TS5,TS6, PS=0,0,0,10,5, rate=10kHz

change to hallc_fadc_ssp config
4126 TS1,TS4, PS1=PS4=0, GEM HVs on after starting the run

========================================================================
20230208

---------------------------------------------------
List of HVs:
board1, 6kV, 8-ch:
Name:_GEM00,_GEM01,_GEM10,_GEM11,__,__,__,_Cer_D
HV:_______0,_____0,_____0,_____0,_0,_0,_0,___950

Name:_SC_D,_SC_A,_SC_B,_X,_LASPD_t,_LASPD_b,_SC_C,_PreSh_L(SDU#1),_PreSh_R(THU#1),_PreSh_T(SDU#2),_Shower_L(SDU#4),_Shower_R(THU#2),_Shower_T(SDU#5),__CerA,__CerB,__CerC
HV:___1150,_1550,_1650,_0,_____800,_____900,_1600,___________1470,___________1740,___________1680,_____________900,____________1030,____________1030,___950,___950,___950

GEM HVs are off all the time.
use hallc_fadc_ssp config

4127-4128 pedestal run, beam off and on, HV off,

4129-4130 pedestal run, beam off, HV off,
4131-4132 pedestal run, beam off, HV on,
4133-4134 pedestal run, beam off, HV on, change Shower HVs (L,R,T) from (900,1030,1030) to (900,1100,1000)

4135 trash, reboot coda
use hallc_fadc config, keep HVs as run 4133, FADC will not record signal below threshold.
4136-4139 pedestal run, no beam, HVs on. 8M events each run
4140 pedestal run, no beam, HVs on. 137M events in this run, 16:00 Feb 8th - 00:05 Feb 9th,
beam was back during from 19:02-19:22 (subrun 5-11) and 22:13-00:05 (subrun 17-37)
data triggers (TS#6 only) are missed with pulse triggers. The timing of TS#6 is out of scope,
therefore these events are all random!

========================================================================
Escorded access 20230210
1. HV channel of 2nd GEM (from upstream to downstream) switched from Board A.1 to A.5.
A.1 is not stable, sometimes it will not work, sometime do.
2. Reconnected SC_A signal cable
3. Reduced the delay cable length of SC_B by 5 feet,
4. We need to have 5 copies of ShowerSum signal. One go to FADC, four go to vmeDSC.
Therefore we need to have an extra FIFO after summing up 3 shower blocks. The copy goes to
FADC is coming right after the summing. All four ShowerSum signals used in vmeDSC are from the extra FIFO.
This FIFO will have ~10ns delay, which will match to CerSum better.
(In the past vmeDSC ch03,ch04,ch05 are all come from summing unit, FADC and vmeDSC ch14 are from the extra FIFO.)

20230210
40uA beam on Carbon target,keep Shower HVs (L,R,T) at (900,1100,1000)
4141 pedestal run, beam on, HV off,
4142 TS4 only, PS4=5, threshold 20mv,
4143 TS4 only, PS4=5, threshold 20mv
4144 TS1 only, PS1=5 threshold 200mv & 20mv
4145 TS1 only, PS1=5 threshold 200mv & 20mv
4146 TS5 only, PS5=5, threshold 10mv
4147 TS2 only, PS2=0, threshold 10mv & 20mv & 10mv
4148 TS2 only, PS2=0, threshold 10mv & 20mv & 10mv
4149 TS4 only, PS4=5, threshold 20mv, change Shower HVs (L,R,T) from (900,1100,1000) to (800,1050,900)

4150-4151 debug GEM, TS5 only, only GEM and SC_B HVs are on, threshold 40mv, PS5=5
4152 TS4 only, GEM and all other HVs are on, threshold 200mv, PS4=0. SC_D DAC value set too low

4153 trash, no raw data file
4154 TS5 only, PS5=3, all HVs are on without GEM, threshold 60mv
4155 TS6 only, PS6=0, all HVs are on without GEM, 15mv 2/15 Cer individual

========================================================================
20230212
15:25 10uA CW beam on He3 target
---------------------------------------------------
List of HVs:
board1, 6kV, 8-ch:
Name:_GEM00,_GEM01,_GEM10,_GEM11,__,__,__,_Cer_D
HV:_______0,_____0,_____0,_____0,_0,_0,_0,___950

Name:_SC_D,_SC_A,_SC_B,_X,_LASPD_t,_LASPD_b,_SC_C,_PreSh_L(SDU#1),_PreSh_R(THU#1),_PreSh_T(SDU#2),_Shower_L(SDU#4),_Shower_R(THU#2),_Shower_T(SDU#5),__CerA,__CerB,__CerC
HV:___1150,_1550,_1650,_0,_____800,_____900,_1600,___________1470,___________1740,___________1680,_____________800,____________1050,_____________900,___950,___950,___950

4156 TS4 only, PS4=3, all HVs are on without GEM, threshold 200mv, rate=470Hz, 66k events
4157 TS4 only, PS4=3, all HVs are on without GEM, threshold 20mv, rate=12.2kHz, 1.7M events, daq saturated
set FADC250_TET_IGNORE_MASK to 0 for slot7 and slot8
4158 TS4 only, PS4=3, all HVs are on without GEM, threshold 20mv(ShowerSum), rate=13.8kHz, 1.7M events
4159 TS1 only, PS1=3, all HVs are on without GEM, threshold 200mv(CerSum, about 12 pe) & 20mv(ShowerSum), rate=13.8kHz, 1.2M events
4160 TS2 only, PS2=0, all HVs are on without GEM, threshold 10mv(SC_D) & 10mv(ShowerSum) & 40mv(SC_B), rate=200Hz, 16k events
4161 TS2 only, PS2=0, all HVs are on without GEM, threshold 6mv(SC_D) & 10mv(ShowerSum) & 6mv(SC_B), rate=13.2kHz, 2.4M events
4162 TS2 only, PS2=0, all HVs are on without GEM, threshold 10mv(SC_D) & 20mv(ShowerSum) & 10mv(SC_B), rate=12.2kHz, 1.1M events
4163 TS3 only, PS3=5, all HVs are on without GEM, threshold 10mv(SC_D) & 50mv(ShowerSum) & 200mv(CerSum, about 12 pe), rate=6kHz, 1.1M events

4164 TS2 only, PS2=1, all HVs are on without GEM, threshold 10mv(SC_D) & 20mv(ShowerSum) & 10mv(SC_B),
beam current went up to 25uA, rate=450Hz (detector saturated if beam current>20uA), 26k events

16:02 Turn aLL HVs off
===4161 good TS2 hreshold: 5mv(SC_D) & 10mv(ShowerSum) & 6mv(SC_B), rate=13.2kHz, 2.4M events===

20230213 reboot NIM crate and vme crate

20230213 threshold
Threshold is applied to the height of the waveform:
TS1: CerSum(50mv or 102 adc unit, about 2.9 pe) & ShowerSum(20mv or 41 adc unit);
TS2: SC_D(6mv or 49 adc unit) & ShowerSum(20mv or 41 adc unit) & SC_B(6mv or 49 adc unit).
TS3: disabled
TS4: ShowerSum(200mv or 405 adc unit)
TS5: SC_B(6mv or 49 adc unit).
TS6: CerA#,CerB#,CerC#,CerD#(15mv or 31 adc unit,about 0.8 pe) for each Cherenkov individual input

set FADC250_TET_IGNORE_MASK to 0 for slot7 and slot8
4165 pedestal run, beam off, HVs on
4166 pedestal run, beam off, HVs on
4167 pedestal run, beam off, HVs off
set FADC250_TET_IGNORE_MASK to 1 for slot7 and slot8
4168 pedestal run, beam off, HVs off
4169 pedestal run, beam off, HVs on

Carbon target
4170 TS4 only, PS4=0, threshold 200mv, change Shower HVs(L,R,T) from (800,1050,900) to (900,1030,1030), daq rate 100Hz, tune beam
4171 TS4 only, PS4=0, threshold 200mv, HVs(L,R,T) stay at (900,1030,1030), daq rate 5kHz, cw beam < 20uA

ToDo:
2uA CW beam, TS2 only, PS2=0, HV scan at the same threshold as 4122:
SC_D(5mv or 41 adc unit) & ShowerSum(5mv or 11 adc unit) & SC_B(6mv or 49 adc unit).
1) HVs(L,R,T) = (900,1030,1030)
2) HVs(L,R,T) = (880,1030,1000)
3) HVs(L,R,T) = (860,1030,975)
4) HVs(L,R,T) = (840,1030,950)
5) HVs(L,R,T) = (820,1030,930)

4172-4189 debug GEM and coda

========================================================================
20230214 Pedestal runs

4190-4191 pedestal runs, beam off, HVs on, hallc_fadc config
4192 pedestal runs, beam off, HVs off, hallc_fadc config
4193 trash, pedestal runs, beam off, HVs off, hallc_fadc_ssp config, ssp fail
4194 pedestal runs, beam off, HVs off, hallc_fadc config
4195 trash, debug ssp
4196-4197 pedestal runs, beam off, HVs off, hallc_fadc_ssp config, With all 4 GEM detectors, daq rate =3.7kHz, saturated
4198-4200 pedestal runs, beam off, HVs on, hallc_fadc_ssp config, With all 4 GEM detectors, daq rate =3.7kHz, saturated

====Only shower and GEM HV on====
40uA CW on dummy
4201 TS6, trash
4202 TS4, PS4=0, threshold: 100mV(ShowerSum) Shower HV (L=700, R=730, T=680 V)
4203 TS4, PS4=0, threshold: 150mV(ShowerSum) Shower HV (L=700, R=730, T=680 V)
4204 TS4, PS4=0, threshold: 150mV(ShowerSum) Shower HV (L=900, R=1030, T=1030 V) change GEM latancy to 59, beam current to high, trash
4205 TS4, PS4=0, threshold: 150mV(ShowerSum) Shower HV (L=900, R=1030, T=1030 V), taken on beam ramp up

===All HV on===
2uA CW beam on dummy tartet
4206 TS4, PS4=0, threshold: 200mv(ShowerSum), rate=2.0kHz Shower HV (L=900, R=1030, T=1030 V)
4207 TS2, PS2=0, threshold: 5mv(ShowerSum), 5mv(SC_D), 6mv(SC_B), rate=3.3kHz Shower HV (L=900, R=1030, T=1030 V)
4208 TS2, PS2=0, threshold: 5mv(ShowerSum), 5mv(SC_D), 6mv(SC_B), rate=2.7kHz Shower HV (L=880, R=1030, T=1010 V)
4209 TS2, PS2=0, threshold: 5mv(ShowerSum), 5mv(SC_D), 6mv(SC_B), rate=2.3kHz Shower HV (L=860, R=1030, T=990 V)
4210 TS2, PS2=0, threshold: 5mv(ShowerSum), 5mv(SC_D), 6mv(SC_B), rate=2.0kHz Shower HV (L=840, R=1030, T=970 V)

Recommented HVs: (L=850, R=1030, T=970 V) ---> MIP at 600
Used in FermiLab beam test: (L=760, R=880, T=860 V)

20230217: test trigger latency
4211-4213 trash, no beam,
4214 TS1, PS1=3, threshold: 50mV (CerSum) & 200mv(ShowerSum), Shower HV (L=760, R=880, T=860 V)
4215 TS2, PS2=0, threshold: 5mv(ShowerSum), 5mv(SC_D), 6mv(SC_B), Shower HV (L=760, R=880, T=860 V)
4216 TS5, PS5=5, threshold: 40mv(SC_B), Shower HV (L=760, R=880, T=860 V)

conclusion:
1) TS5: SC_B signal is 200ns earlier than TS5
2) TS4: Shower_L,Shower_R,Shower_T are 200ns earlier than TS4, ShowerSum is 6ns later than Shower_L
3) TS1: only one peak, but showers have a wide range of 12~14 windows
4) TS2: only one peak, but most detectors have a wide range of 14-16 windows
5) SC_D, SC_B, CerSum have 44ns,38ns,44ns cable delay. SC_B is 6ns later than SC_D before delay.
After these delay, SC_B is 16ns later than ShowerSum, CerSum is 4ns later than ShowerSum.
We need to reduce the delay cable length for SC_B, SC_D and CerSum.
6) TS6 is 352ns (or 88 windows) behind its input signals

Trigger signal at Slot7_ch07 is 9 windows eariler than the trigger signal at slot8_ch15. Because Slot7_ch07 signal
is only 1 window later than that of Slot8_ch15, we should set FADC250_W_OFFSET 10 windows 40ns eariler in order to
match these 2 slots in timing. Therefore I changed slot7 from "FADC250_W_OFFSET 1060" to "FADC250_W_OFFSET 1100".
I also changed slot7 FADC250_ALLCH_DELAY from 160ns to 80ns.
FADC250_ALLCH_DELAY 80 80 80 80 80 80 80 0 80 80 80 80 80 80 80 80

4217 will be the first run after FADC slo7 config change in FADC250_ALLCH_DELAY and FADC250_W_OFFSET.

========================================================================
4217 TS4 good data, need to verify if Shower HV = (760,880,860)?
4218 TS2 good data, need to verify if Shower HV = (760,880,860)?

check TS6 latency, trash run are removed
4219-4221 Trash
4222-4225 TS6, only slot8 is useful, Shower HV = (760,880,860), peak location is 5 windows earilier than TS4
4226-4238 Trash
4239-4240 TS6, only slot8 is useful, Shower HV = (850,1030,970), peak location is 5 windows earilier than TS4
4241-4242 Trash
4243-4247 TS6, only slot8 is useful, Shower HV = (850,1030,970), peak location is the same as TS4
4248-4250 TS4, Shower HV = (850,1030,970),
4251 TS2, Shower HV = (850,1030,970),
4252-4260 Trash, FADC need to reboot
4261 TS4, only Shower HVs are on. Shower HV = (850,1030,970), TS4 delay = 40 windows, first 100k are pulse triggers
4262 TS1 and TS2, all HVs are on without GEM. Shower HV = (850,1030,970), TS1 TS2 delay = 4,7 windows
4263 TS1 and TS2, all HVs are on without GEM. Shower HV = (850,1030,970), TS1 TS2 delay = 4,7 windows
//////////////////////////////////////////////////////////////////////////////////////
no TS delay
4264 TS4, PS4=3, thredhold=200mv(ShowerSum), rate=1.5kHz
4265 TS4, PS4=3, thredhold=40mv(ShowerSum), rate=12.8kHz, daq Saturated
4266 TS4, PS4=3, thredhold=40mv(ShowerSum), rate=12.8kHz, daq Saturated
4267 TS4, PS4=5, thredhold=40mv(ShowerSum), rate=12.8kHz, daq Saturated
4268 TS4, PS4=10, thredhold=40mv(ShowerSum), rate=7.8kHz,
4269 TS4, PS4=10, thredhold=40mv(ShowerSum), rate=7.8kHz,
4270 TS4, PS4=10, thredhold=40mv(ShowerSum), rate=7.8kHz,
4271 TS4, PS4=10, thredhold=40mv(ShowerSum), rate=7.8kHz,
4272 TS1, PS1=5, thredhold=50mv(CerSum),20mv(ShowerSum), rate=12.8kHz, daq Saturated
4273 trash
4274 TS1, PS1=10, thredhold=50mv(CerSum),20mv(ShowerSum), rate=7.8kHz
///////////////////////////////////////////////////////////////////////////////////////

How to determine TS6 latency:
basic knowledge:
A) TS6 (FADC trigger) is ~88 windows behind FADC input signal (viewed in DiagGui).
B) TS4 is about 49 windows behind the input signal of logic unit (viewed in data).

1) Use TS4 (NIM trigger, trigger by slot8_ch14), plug the trigger signal(L1A) into FADC slot7 and slot8.
Modify FADC250_W_OFFSET and FADC250_ALLCH_DELAY of FADC config file such that the peaks of
these 2 signals stay at the same time in slot7 and slot8. The following is my current setting.
Trigger signals have ZERO delay.

slot7: FADC250_W_OFFSET 1100
slot7: FADC250_ALLCH_DELAY 80 80 80 80 80 80 80 0 80 80 80 80 80 80 80 80
slot8: FADC250_W_OFFSET 1060
slot8: FADC250_ALLCH_DELAY 40 80 40 40 80 80 80 40 40 40 40 40 40 40 40 0
Here is the peak location of each channel.
double kSlot7PeakPos_416 = { 44.5, 43.5, 44.5, 44.5, 43.5, 43.5, 43.5, 78.5, 43.5, 43.5, 43.5, 43.5, 43.5, 43.5, 43.5, 43.5 };
double kSlot8PeakPos_416 = { 40.5, 35.5, 41.5, 39.5, 38.5, 37.5, 31.5, 35.5, 35.5, 36.5, 37.5, 37.5, 37.5, 37.5, 38.5, 77.5 };
Note that the trigger signal of slot7_ch07 is 6ns late than that of slot8_ch14 because it gord thrught an extra FIFO.

2) From the above peak location, we can see that TS6_peak = 43.5 + 94 = 137.5 (windows). To align TS4 to TS6, we have to
delay TS4 by 137.5 - 77.5 = 60 windows. In my case, I change the readout list, delay TS4 by 54 windows.
We fisrtly change FADC250_W_OFFSET to view the peak location of these 2 trigger signals.
To keep slot7_ch07 at 78.5, slot7_W_offset should be 1100 - 60x4 = 860.
To keep slot8_ch15 at 77.5, slot8_W_offset should be 1060 - 60x4 = 820.
Since FADC trigger is 88 windows later than its inputs, it is impossible to view inputs and trigger within a 400ns windows.
Therefore do not bother to change slot7 delay. But we can add 60 windows (240ns) delay to slot8 such that the peaks of all
channels stay the same location.

slot7: FADC250_W_OFFSET 860
slot7: FADC250_ALLCH_DELAY 80 80 80 80 80 80 80 0 80 80 80 80 80 80 80 80
slot8: FADC250_W_OFFSET 820
slot8: FADC250_ALLCH_DELAY 256 296 256 256 296 296 296 256 256 256 256 256 256 256 256 0

Take a run using TS6 trigger with the above changes to determine the peak location for 2 trigger signals.

========================================================================
20230219 run list
All HVs are on without GEM. Shower HV = (850,1030,970).
40uA CW beam on Li-7 target.

---------------------------------------------------
List of HVs:
board1, 6kV, 8-ch:
Name:_GEM00,__,__,_GEM01,_GEM10,_GEM11,__,_Cer_D
HV:_______0,_0,_0,_____0,_____0,_____0,_0,___950

Name:_SC_D,_SC_A,_SC_B,_X,_LASPD_t,_LASPD_b,_SC_C,_PreSh_L(SDU#1),_PreSh_R(THU#1),_PreSh_T(SDU#2),_Shower_L(SDU#4),_Shower_R(THU#2),_Shower_T(SDU#5),__CerA,__CerB,__CerC
HV:___1150,_1550,_1650,_0,_____800,_____900,_1600,___________1470,___________1740,___________1680,_____________850,____________1030,_____________970,___950,___950,___950

4275 TS4, PS=10, Threshold: 40mV(ShowerSum), rate=8.2kHz, Shower HV = (850,1030,970)
4276 TS2, PS=0, Threshold: 5mV(SC_D) & 5mV(ShowerSum) & 6mV(SC_B), Shower HV = (850,1030,970),
rate=12.5kHz then detectors got saturated. 170k triggers
4277 TS2, PS=0, Threshold: 5mV(SC_D) & 5mV(ShowerSum) & 6mV(SC_B), Shower HV = (850,1030,970),
rate=12.5kHz then detectors got saturated. 988k triggers
4278 TS2, PS=0, Threshold: 5mV(SC_D) & 5mV(ShowerSum) & 6mV(SC_B), Shower HV = (850,1030,970),
rate=12.5kHz then detectors got saturated. 508k trigger
4279 TS2, PS=0, Threshold: 5mV(SC_D) & 5mV(ShowerSum) & 6mV(SC_B), Shower HV = (840,1030,970),
rate=12.5kHz then detectors got saturated. 856k triggers
4280 TS2, PS=0, Threshold: 5mV(SC_D) & 5mV(ShowerSum) & 6mV(SC_B), Shower HV = (840,1030,970),
rate=12.5kHz then detectors got saturated. 526k triggers

4281 trash
4282 pedestal. Beam off, HV off.
4283 pedestal. Beam off, HV off.
4284 pedestal. Beam off, HV on. Shower HV = (840,1030,970),
4285 pedestal. Beam off, HV on. Shower HV = (850,1030,970).

==================== End of 7 deg ====================================

For 18deg data set, use this link: https://redmine.jlab.org/projects/solid-ecal/wiki/Hallc_beamtest_runlist_18deg?parent=2022_JLab_Hall_C_Beam_test

Updated by Jixie Zhang 12 months ago · 3 revisions